# An Incomplete Loop: Deductive, Inductive, and Abductive Learning in Large Language Models

### v2

@@ Appendix H Vocabulary Induction Accuracy for Colours Domain
gpt-3.5-turbo lug verbal_conf 0.604859335038363 -0.0370824643671235 0.300350893877687 0.52561406
gpt-3.5-turbo dax verbal_conf 0.233072916666666 0.0669078409779957 0.0638449077681366 0.29564626
gpt-3.5-turbo wif verbal_conf 0.329268292682926 0.0360864838621108 0.646415727402187 0.82271093
gpt-3.5-turbo zup verbal_conf 0.61 0.154599744147798 0.0288284171902121 0.20179892
@@
gpt-4-turbo lug verbal_conf 0.875763747454175 0.0439535251603629 0.168735841765082 3.58e-1
gpt-4-turbo dax verbal_conf 0.921487603305785 -0.0418339829001646 0.193444842357041 3.58e-1
gpt-4-turbo wif verbal_conf 0.865853658536585 0.0447789368805147 0.569118855838926 7.32e-1
gpt-4-turbo zup verbal_conf 0.87 0.0579317588377811 0.415170703498878 5.75e-1
gpt-4-turbo bluf verbal_conf 0.153846153846153 0.130289326661761 0.255554806693343 3.83e-1
gpt-4-turbo walm verbal_conf 0.171428571428571 0.0489476613672 0.687394764355678 7.55e-1
gpt-3.5-turbo lug p_data 0.28772378516624 0.0591580864392629 0.0983076746979401 0.29564626
gpt-3.5-turbo dax p_data 0.10546875 0.00996669750420649 0.782730731634431 0.84294079
gpt-3.5-turbo wif p_data 0.195121951219512 0.24388397206115 0.00164950240178991 0.02309303
gpt-3.5-turbo zup p_data 0.4 -0.114775306453679 0.105587948530317 0.29564626
gpt-3.5-turbo bluf p_data 0.0769230769230769 -0.1673596703488 0.143038746170574 0.33375707
@@
gpt-4-turbo lug p_data 0.731160896130346 -0.0803208788270973 0.0118067271484766 5.31e-2
gpt-4-turbo dax p_data 0.733471074380165 -0.0411584340901287 0.200743028038944 3.58e-1
gpt-4-turbo wif p_data 0.768292682926829 -0.0255095864281423 0.745759257358691 7.55e-1
gpt-4-turbo zup p_data 0.76 0.0221788843765235 0.755246389745868 7.55e-1
gpt-4-turbo bluf p_data 0.102564102564102 -0.140859042454752 0.218662613923779 3.58e-1
gpt-4-turbo walm p_data 0.142857142857142 -0.263591223386182 0.0274677998646683 9.89e-2
gpt-3.5-turbo lug p_answer 0.436061381074168 -0.0128717673177528 0.719304414221268 0.83918848
gpt-3.5-turbo dax p_answer 0.190104166666666 0.00399825404003988 0.9119154429019 0.91191544
gpt-3.5-turbo wif p_answer 0.353658536585365 -0.0743944887571608 0.343771340895407 0.53475542
gpt-3.5-turbo zup p_answer 0.59 -0.0879883939992145 0.21536453378319 0.43072907
gpt-3.5-turbo bluf p_answer 0.0512820512820512 -0.0551410966570355 0.631586163340137 0.82271093
@@
gpt-4-turbo lug p_answer 0.70468431771894 -0.0661341939950676 0.0382590601605542 1.15e-1
gpt-4-turbo dax p_answer 0.712809917355371 0.188125007589795 3.67e-9 6.61e-8
gpt-4-turbo wif p_answer 0.817073170731707 -0.249727938065423 0.00126030696185871 1.13e-2
gpt-4-turbo zup p_answer 0.68 -0.142141439177509 0.0446656422362514 1.15e-1
gpt-4-turbo bluf p_answer 0.23076923076923 -0.294147299400922 0.00894893196851883 5.31e-2
gpt-4-turbo walm p_answer 0.0571428571428571 -0.0578005412233204 0.634584373040489 7.55e-1
@@ Appendix I MTOB Experimental Settings
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
Hypothesis proposal system prompt (Grammar)
@@
For proposing grammar feature hypotheses based on in-context examples
Hypothesis proposal system prompt (Vocab)
@@
For proposing vocabulary mappings based on in-context examples
Instruction following system prompt
@@
For applying a proposed hypothesis or ground-truth hypothesis to the input
Few-shot examples prompt
For reasoning with in-context examples only
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Now write the translation. English: {query input} Kalamang translation:
Prompt with ground-truth hypothesis
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the translation of a word.
The following sentences contain the word ‘{word}’ in English. Examples: English sentence: {eng sentence} Kalamang translation: {kgv sentence} ... What is the Kalamang translation of the word ‘{word}’? Write your answer like this: {English word} -> {Kalamang translation}.
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{hypothesis translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {hypothesis grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
@@
Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Let’s think step by step about what the translation could be. Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
These are examples of the translation of the word {word}. Examples: {example translation pairs} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of sentences that contain the word {word}: This is the translation of the word: {model’s word hypothesis} Examples: {example translation pairs}
@@
Source
Reference
Translation (few-shot)
Translation (instruction-induction:verbalized confidence)
Translation (true-instruction)
Who asked you?
Namana kat gerket?
Naman mua kiemara kowam?
Naman era?
Naman mua
@@
The rope is wrapped around it.
Karek mat kokarap.
Kainasu esun kiwat semda bo dalangdalang.
Sar tok esun et dalangdalang na.
Karek mambon guanggarien ma he me.
Before, they said Japan and The Netherlands were at war.
Wiseme mu toni Japangbon Beladarbon nausair.
Wiseme mu Jafaris emun Nabil emun kaden kiesun ning.
Wise me mu salai Japan an Sar Beladargara din jambat mara.
Tikninda, termus siram panci tubak Beladar tamatko pang.
Mu bo Tamisengga som dande yecie.
They are going (on a return trip) to bury someone in Antalisa.
They go to Tamisengga and return.
They go to Antalisa, someone died, and they return.
They will return to Tamisengga until they bury the person.
Suo me kene arkon metko.
At Suo is a kene tree.
I washed my body in the river.
Back me gives birth canoe.
Wash your face near the Narkon tree.
An istrat kahendengoa marmar ba temun.
I’m walking on a long wide road.
I put the long stick on the ground.
I have a long drying rack but I am walking.
I walked far down the street to my uncle.

### v3

@@ Appendix H Vocabulary Induction Accuracy for Colours Domain
gpt-3.5-turbo lug verbal_conf 0.60 -0.037 0.30 0.53
gpt-3.5-turbo dax verbal_conf 0.23 0.067 0.064 0.30
gpt-3.5-turbo wif verbal_conf 0.33 0.036 0.65 0.82
gpt-3.5-turbo zup verbal_conf 0.61 0.15 0.029 0.20
@@
gpt-4-turbo lug verbal_conf 0.88 0.044 0.17 3.58e-1
gpt-4-turbo dax verbal_conf 0.92 -0.042 0.19 3.58e-1
gpt-4-turbo wif verbal_conf 0.86 0.0445 0.57 7.32e-1
gpt-4-turbo zup verbal_conf 0.87 0.058 0.42 5.75e-1
gpt-4-turbo bluf verbal_conf 0.15 0.13 0.26 3.83e-1
gpt-4-turbo walm verbal_conf 0.17 0.049 0.69 7.55e-1
gpt-3.5-turbo lug p_data 0.29 0.059 0.098 0.30
gpt-3.5-turbo dax p_data 0.10 0.010 0.78 0.84
gpt-3.5-turbo wif p_data 0.20 0.24 0.0016 0.023
gpt-3.5-turbo zup p_data 0.4 -0.11 0.10 0.30
gpt-3.5-turbo bluf p_data 0.077 -0.17 0.14 0.33
@@
gpt-4-turbo lug p_data 0.73 -0.080 0.012 5.31e-2
gpt-4-turbo dax p_data 0.73 -0.041 0.20 3.58e-1
gpt-4-turbo wif p_data 0.77 -0.026 0.74 7.55e-1
gpt-4-turbo zup p_data 0.76 0.022 0.76 7.55e-1
gpt-4-turbo bluf p_data 0.10 -0.14 0.22 3.58e-1
gpt-4-turbo walm p_data 0.14 -0.26 0.027 9.89e-2
gpt-3.5-turbo lug p_answer 0.44 -0.013 0.72 0.84
gpt-3.5-turbo dax p_answer 0.19 0.0040 0.91 0.91
gpt-3.5-turbo wif p_answer 0.35 -0.074 0.34 0.53
gpt-3.5-turbo zup p_answer 0.59 -0.088 0.21 0.43
gpt-3.5-turbo bluf p_answer 0.051 -0.055 0.63 0.82
@@
gpt-4-turbo lug p_answer 0.70 -0.066 0.038 1.15e-1
gpt-4-turbo dax p_answer 0.71 0.19 3.67e-9 6.61e-8
gpt-4-turbo wif p_answer 0.82 -0.25 0.0013 1.13e-2
gpt-4-turbo zup p_answer 0.68 -0.14 0.045 1.15e-1
gpt-4-turbo bluf p_answer 0.23 -0.29 0.0089 5.31e-2
gpt-4-turbo walm p_answer 0.057 -0.058 0.63 7.55e-1
@@ Appendix I MTOB Experimental Settings
Prompt Type Usage Prompt Text
Base system prompt For reasoning with in-context examples ‘‘’’
@@
Hypothesis proposal system prompt (Grammar) For proposing grammar feature hypotheses based on in-context examples ‘‘’’
@@
Hypothesis proposal system prompt (Vocab) For proposing vocabulary mappings based on in-context examples ‘‘’’
@@
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input ‘‘’’
@@
Few-shot examples prompt For reasoning with in-context examples only Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Now write the translation. English: {query input} Kalamang translation:
@@
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the translation of a word. The following sentences contain the word ‘{word}’ in English. Examples: English sentence: {eng sentence} Kalamang translation: {kgv sentence} ... What is the Kalamang translation of the word ‘{word}’? Write your answer like this: {English word} -> {Kalamang translation}.
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is one of the closest entries to ‘{word}‘ in the Kalamang=English bilingual dictionary: English word or phrase: ‘{word}‘ Kalamang translation: ‘{hypothesis translation}’ <Repeated for all words> To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> To help with the translation, here’s a grammar sketch of Kalamang: {hypothesis grammar sketch} Now write the translation. English: {query input} Kalamang translation:
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Human: Kalamang is a language spoken on the Karas Islands in West Papua. Translate the following sentence from English to Kalamang: {query input} To help with the translation, here is a translated sentence with words similar to ‘{word}’ in a list of translated Kalamang-English reference sentences: English sentence: {eng sentence} Kalamang translation: {kgv sentence} <Repeated for all words> Let’s think step by step about what the translation could be. Now write the translation. English: {query input} Kalamang translation:
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses These are examples of the translation of the word {word}. Examples: {example translation pairs} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of sentences that contain the word {word}: This is the translation of the word: {model’s word hypothesis} Examples: {example translation pairs}
@@
Source Reference Translation (few-shot) Translation (instruction-induction:verbalized confidence) Translation (true-instruction)
Who asked you? Namana kat gerket? Naman mua kiemara kowam? Naman era? Naman mua
@@
The rope is wrapped around it. Karek mat kokarap. Kainasu esun kiwat semda bo dalangdalang. Sar tok esun et dalangdalang na. Karek mambon guanggarien ma he me.
Before, they said Japan and The Netherlands were at war. Wiseme mu toni Japangbon Beladarbon nausair. Wiseme mu Jafaris emun Nabil emun kaden kiesun ning. Wise me mu salai Japan an Sar Beladargara din jambat mara. Tikninda, termus siram panci tubak Beladar tamatko pang.
Mu bo Tamisengga som dande yecie. They are going (on a return trip) to bury someone in Antalisa. They go to Tamisengga and return. They go to Antalisa, someone died, and they return. They will return to Tamisengga until they bury the person.
Suo me kene arkon metko. At Suo is a kene tree. I washed my body in the river. Back me gives birth canoe. Wash your face near the Narkon tree.
An istrat kahendengoa marmar ba temun. I’m walking on a long wide road. I put the long stick on the ground. I have a long drying rack but I am walking. I walked far down the street to my uncle.